OpenStudy (anonymous):

Please help!! Factor Completely: -3x + 8

mathslover (mathslover):

Can I write 8 as 5 + 3 ?

OpenStudy (anonymous):

uh... however you think

OpenStudy (anonymous):

no. well, i dunno. Its short for Ashley Is Epic =P

mathslover (mathslover):

:) \(-3x + 3 + 5\) = \(-3x + 8\) as 8 = 5+ 3, so I can put (5+3) on the place of 8 . \(-3x + 3 + 5 = -3(x-1) + 5 \) \(-3x + 3 + 5 = -3(x-1) + 3 + 2 = -3(x-1 - 1) + 2 = -3(x-2) + 2\) So, you can write -3x+8 as -3(x-2) + 2.

OpenStudy (anonymous):

oh okay. Thank you!